2) Burke Company received an invoice dated November 18 for a shipment of goods received November 21. The invoice was for $9255.00 less 40%, 13.375% with terms 5/10, 3/20, n/45. How much must Burke pay on December 2 to reduce its debt a) by $2000.00? b) to $2000.00? c) if Burke Company makes a partial payment of $4000 on November 25, how much should be credited to the account?
Answer #1
Invoice amt. = 9255
Net amount due after successive discounts
9255*(1-40%)*(1-13.375%)=
4810.29
Payment by Burke on Dec 2
to reduce its debt
Dec 2 falls after 10 days but within 20 days
to which 3% discount applies
So,to reduce its debt BY 2000
a) by $ 2000
Payment to be made will be
2000*(1-3%)=
1940
b) reduce its debt TO $2000.00
ie. Payment to be made will be
(4810.29-2000)*(1-3%)=
2725.98
c) if Burke Company makes a partial payment of $4000 on November 25, amount credited to the account:
Nov. 25 falls within 10 days,
to which 5% discount applies
so, when a payment of 4000 is made,
amt. credited will be
4000/(1-5%)*1=
4210.53